# MetaGFN: Exploring Distant Modes with Adapted Metadynamics for Continuous GFlowNets

### v1

@@ Abstract
Generative Flow Networks (GFlowNets) are a class of generative models that sample objects in proportion to a specified reward function through a learned policy. They can be trained either on-policy or off-policy, needing a balance between exploration and exploitation for fast convergence to a target distribution. While exploration strategies for discrete GFlowNets have been studied, exploration in the continuous case remains to be investigated, despite the potential for novel exploration algorithms due to the local connectedness of continuous domains. Here, we introduce Adapted Metadynamics, a variant of metadynamics that can be applied to arbitrary black-box reward functions on continuous domains. We use Adapted Metadynamics as an exploration strategy for continuous GFlowNets. We show three continuous domains where the resulting algorithm, MetaGFN, accelerates convergence to the target distribution and discovers more distant reward modes than previous off-policy exploration strategies used for GFlowNets.
@@ 1 Introduction
Generative Flow Networks (GFlowNets) are a type of generative model that samples from a discrete space $\mathcal{X}$ by sequentially constructing objects via actions taken from a learned policy $P_{F}$ Bengio et al. (2021a). The policy $P_{F}(s,s^{\prime})$ specifies the probability of transitioning from some state $s$ to some other state $s^{\prime}$. The policy is parameterised and trained so that, at convergence, the probability of sampling an object $x\in\mathcal{X}$ is proportional to a specified reward function $R(x)$. GFlowNets offer advantages over more traditional sampling methods, such as Markov chain Monte Carlo (MCMC), by learning an amortised sampler, capable of single-shot generation of samples from the desired distribution. Since GFlowNets learn a parametric policy, they can generalise across states, resulting in higher performance across various tasks Bengio et al. (2021a); Malkin et al. (2022); Zhang et al. (2022); Jain et al. (2022); Deleu et al. (2022); Jain et al. (2023); Hu et al. (2023); Zhang et al. (2023); Shen et al. (2023b) and applications to conditioned molecule generation Shen et al. (2023b), maximum likelihood estimation in discrete latent variable models Hu et al. (2023), structure learning of Bayesian networks Deleu et al. (2022), scheduling computational operations Zhang et al. (2023), and discovering reticular materials for carbon capture Cipcigan et al. (2023).
@@
Although originally conceived for discrete state spaces, GFlowNets have been extended to more general state spaces, such as entirely continuous spaces, or spaces that are hybrid discrete-continuous Lahlou et al. (2023). In the continuous setting, given the current state, the policy specifies a continuous probability distribution over subsequent states, and the probability density over states $x\in\mathcal{X}$ sampled with the policy is proportional to a reward density function $r(x)$. The continuous domain unlocks more applications for GFlowNets, such as molecular conformation sampling Volokhova et al. (2023) and continuous control problems Luo et al. (2024).
@@
GFlowNets are trained like reinforcement learning agents. Trajectories of states are generated either on-policy or off-policy, with the terminating state $x\in\mathcal{X}$ providing a reward signal for informing a gradient step on the policy parameters. GFlowNets therefore suffer from the same training pitfalls as reinforcement learning. One such issue is slow temporal credit assignment, which has thus far been addressed by designing more effective loss functions, such as detailed balance Bengio et al. (2021b), trajectory balance Malkin et al. (2022) and sub-trajectory balance Madan et al. (2022).
@@
Besides loss functions, another aspect of GFlowNet training is the exploration strategy for acquiring training samples. Exclusively on-policy learning is generally inadequate as it leads to inefficient exploration of new modes. More successful strategies therefore rely on off-policy exploration. For the discrete setting, numerous exploration strategies have been proposed including $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) Pan et al. (2022), Thompson sampling Rector-Brooks et al. (2023) and Local Search GFlowNets Kim et al. (2024). While these approaches can be generalised to the continuous domain, there is no literature benchmarking their effectiveness in this setting.
Sampling in the continuous setting is a common occurrence in various domains such as molecular modelling Hawkins (2017); Yang et al. (2019) and Bayesian inference Shahriari et al. (2016). The local connectedness of a continuous domain allows for novel exploration strategies that are not directly applicable in the discrete setting. In this work, we create MetaGFN, an exploration algorithm for continuous GFlowNets inspired by metadynamics, an enhanced sampling method widely used for molecular modelling Laio and Parrinello (2002). The main contributions of this work are:
@@
  • Presenting MetaGFN, an algorithm created by adapting metadynamics to black box rewards and continuous GFlowNets;
@@
  • Proving that the proposed Adapted Metadynamics is consistent and reduces to standard metadynamics in a limit;
@@
  • Showing empirically that MetaGFN outperforms existing GFlowNets exploration strategies.
@@
The rest of the paper is as follows. In Section 2 we review the theory of discrete and continuous GFlowNets. We present Adapted Metadynamics and MetaGFN in Section 3. In Section 4, we evaluate MetaGFN against other exploration strategies, showing that MetaGFN outperforms existing exploration strategies in two continuous environments. We finish with limitations and conclusions in Sections 5 and 6. Code for MetaGFN is available at [link in camera-ready].
@@ 2.1 Discrete GFlowNets
In a GFlowNet, the network refers to a directed acyclic graph (DAG), denoted as $G=(\mathcal{S},\mathcal{A})$. Nodes represent states $s\in\mathcal{S}$, and edges represent actions $s\rightarrow s^{\prime}\in\mathcal{A}$ denoting one-way transitions between states. The DAG has two distinguishable states: a unique source state $s_{0}$, that has no incoming edges, and a unique sink state $\perp$, that has no outgoing edges. The set of states, $\mathcal{X}\subset\mathcal{S}$, that are directly connected to the sink state are known as terminating states. GFlowNets learn forward transition probabilities, known as a forward policy $P_{F}(s^{\prime}|s)$, along the edges of the DAG so that the resulting marginal distribution over the terminal states complete trajectories (the terminal distribution) , denoted as $P^{\perp}(x)$, is proportional to a given reward function $R\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$. GFlowNets also introduce additional learnable objects, such as a backward policy $P_{B}(s|s^{\prime})$, which is a distribution over the parents of any state of the DAG, to create losses that train the forward policy. Objective functions for GFlowNets include flow matching (FM), detailed balance (DB), trajectory balance (TB) and subtrajectory balance (STB) Bengio et al. (2021a, b); Malkin et al. (2022); Madan et al. (2022). During training, the parameters of the flow objects are updated with stochastic gradients of the objective function applied to batches of trajectories. These trajectory batches can be obtained either directly from the current forward policy or from an alternative algorithm that encourages exploration. These approaches are known as on-policy and off-policy training respectively.
@@ 2.2 Continuous GFlowNets
Continuous GFlowNets extend the generative problem to continuous spaces Lahlou et al. (2023), where the analogous quantity to the DAG is a measurable pointed graph (MPG) Nummelin (1984). MPGs can model continuous spaces (e.g., Euclidean space, spheres, tori), as well as hybrid spaces, with a mix of discrete and continuous components, as often encountered in robotics, finance, and biology Neunert et al. (2020); Swiler et al. (2012); Bortolussi and Policriti (2008).
@@
    A fully-explorable space: The number of steps required to be able to reach any measurable $B\in\Sigma$ from the source state with the forward reference kernel is bounded.
@@
The reward measure is a positive and finite measure $R$ over the terminating states $\mathcal{X}$, we denote the density of this reward measure as $r(x)$, for $x\in\mathcal{X}$. A flow $F$ is said to satisfy the reward-matching conditions if
@@
If a flow satisfies the reward-matching conditions and trajectories are recursively sampled from the Markov kernel $P_{F}$ starting at $s_{0}$, the resulting measure over terminating states, $P^{\perp}(B)$, is proportional to the reward: $P^{\perp}(B)=\frac{R(B)}{R(\mathcal{X})}$ for any $B$ in the $\sigma$-algebra of terminating states Lahlou et al. (2023).
@@ 2.3 Exploration strategies for GFlowNets
GFlowNets can reliably learn using off-policy trajectories, a key advantage over hierarchical variational models Malkin et al. (2023). For optimal training, it is common to use a replay buffer and alternate between on-policy and off-policy (exploration) batches Shen et al. (2023a). Exploration strategies for discrete GFlowNets include $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) Pan et al. (2022), Local Search GFlowNets Kim et al. (2024), and Thompson sampling (TS), which outperforms the others in grid and bit sequence domains Rector-Brooks et al. (2023). TS aims to bias exploration in regions where there is high uncertainty. When the forward policy is parameterised as an MLP, this is achieved using an ensemble of $K\in\mathbb{Z}^{+}$ policy heads with a common network torso. During training, an ensemble member is randomly sampled and used to generate a trajectory $\tau$. In a training batch, each ensemble member is included with probability $p$ and parameters are updated by taking a gradient step on the total loss of $\tau$ over all included members.
No comparative literature exists on exploration strategies for continuous GFlowNets, but many methods can be adapted, and we do so here. For example, for TS in the continuous setting, policy heads parameterise forward policy functions instead of log probabilities. Another strategy unique to the continuous setting is what we call noisy exploration. This strategy involves introducing an additive noise parameter, denoted as $\bar{\sigma}$, to the variance parameters in the forward policy distribution, where the value of $\bar{\sigma}$ is scheduled to gradually decrease to zero over the course of training.
2.4 Metadynamics
Molecular dynamics (MD) uses Langevin dynamics (LD) Pavliotis (2014), a stochastic differential equation modelling particle motion with friction and random fluctuations, to simulate atomic trajectories that ergodically sample a molecule’s Gibbs measure, $\rho_{\beta}(x)\propto e^{-\beta V(x)}$. Here, $x$ and $p$ are atomic positions and momenta, $V(x)$ is the molecular potential, and $\beta$ is thermodynamic beta. [^1] We review Langevin dynamics in Appendix B. If the potential $V(x)$ has multiple local minima, then unbiased LD can get trapped in these minima, which can lead to inefficient sampling. Metadynamics is an algorithm that enhances sampling by regularly depositing repulsive Gaussian bias potentials at the centre of an evolving LD trajectory Laio and Parrinello (2002). At time $t>0$, metadynamics modifies LD such that the conservative component of the LD force is given by the negative gradient of the total potential, i.e. $-\nabla V_{\text{total},t}=-\nabla(V+V_{\text{bias},t})$, where $V_{\text{bias},t}$ is the cumulative bias at time $t$. Bias potentials are typically chosen to only vary in the direction of user-specified, low-dimensional collective variables (CVs), $z(x):\mathcal{X}\rightarrow\mathcal{Z}$, mapping from the original space $\mathcal{X}$ to the CV space $\mathcal{Z}$. For biomolecules, typical CVs include protein backbone angles or distances between charge centres; these are quantities that drive and parameterise rare-event transitions. As the bias potential progressively fills up the potential landscape, energetic barriers are reduced, thus accelerating exploration, eventually ensuring uniform diffusion in CV space. In practice, the bias is defined on a regular grid, which limits CV space dimensionality to 5-10 due to exponential memory costs. Thus, identifying good CVs is crucial for effective metadynamics simulations in applications such as drug discovery, chemistry and materials science De Vivo et al. (2016); Laio and Gervasio (2008). In this work, we adapt the original metadynamics algorithm discussed above. The method has also seen numerous extensions. For a detailed review, see Bussi and Laio (2020) and the references therein Bussi and Laio (2020).
@@ 3 MetaGFN: Adapted Metadynamics for GFlowNets
Training GFlowNets in high-dimensional continuous spaces requires exploration, especially if valuable reward peaks are separated by large regions of low reward. In some tasks, prior knowledge of the principal manifold directions along which the reward measure varies can reduce the effective dimension required for exploration. This is where exploration algorithms that guarantee uniform sampling in that low-dimensional manifold, such as metadynamics, become most effective. With this intuition in mind, we adapt metadynamics to the black-box reward setting of continuous GFlowNets.
@@
Kernel density potential Metadynamics force computations require the gradient of the total potential, where $-\nabla V_{\text{total},t}=-\nabla(V+V_{\text{bias},t})$. Using the above assumptions, we have $\nabla V=-\nabla r(x)/(\beta^{\prime}r(x))$. However, $r(x)$ is often a computationally expensive black-box function, and its gradient, $\nabla r(x)$, is unknown. While finite differences can estimate $\nabla r(x)$ for smooth, low-dimensional reward distributions, this approach is impractical in high-dimensional spaces. We avoid finite difference calculations by computing a kernel density estimate (KDE) of $V$. In particular, we assume collective variables $z(x)=(z_{1}(x),\dots,z_{k}(x))$, and compute both the KDE potential and bias potentials in the CV space. Here, each $z_{i}$ is a one-dimensional coordinate, and $z\vcentcolon\mathcal{X}\rightarrow\mathcal{Z}$, where $\mathcal{Z}$ is $k$-dimensional. As the potentials are stored on a low-dimensional grid, gradient computations are guaranteed to be cheap relative to the cost of evaluating $r(x)$. We use $\hat{V}(z)$ to denote the KDE estimate in CV space.
To compute $\hat{V}(z)$, we maintain two separate KDEs: $\hat{N}(z)$ for the histogram of metadynamics states and $\hat{R}(z)$ for cumulative rewards. We update these KDE estimates on the fly at the same time the bias potential is updated. If $\mathcal{Z}\cong\mathbb{R}^{k}$, we use Gaussian kernels with update rules: [^2] If $\mathcal{Z}\cong\mathbb{T}^{k}$ ($k$-torus), we use von Mises distributions instead of Gaussians.
$\hat{N}(z)_{t}=\hat{N}(z)_{t-1}+\exp\left(-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-z_{t,i}}{\sigma_{i}}\right|^{2}\right),$
$\hat{R}(z)_{t}=\hat{R}(z)_{t-1}+r(x_{t})\exp\left(-\frac{1}{2}\sum_{i=1}^{k}\left|\frac{z_{i}-z_{t,i}}{\sigma_{i}}\right|^{2}\right),$
where $\sigma=(\sigma_{1},\ldots,\sigma_{k})\in\mathbb{R}^{k}$ is the kernel width, $x_{t}$ is the latest metadynamics sample and $z_{t,i}$ is the corresponding $i^{th}$ CV coordinate of $x_{t}$. The KDE potential $\hat{V}(x)$ is then computed as:
$\hat{V}(x)=-\frac{1}{\beta^{\prime}}\log\left(\frac{\hat{R}(x)}{\hat{N}(x)+\epsilon}+\epsilon\right),$
where $\epsilon>0$. We found empirically that $\epsilon$ ensured numerical stability by preventing division by zero and bounding the potential above by $\log\left(1/\epsilon\right)/\beta^{\prime}$. Taking the ratio of a reward and frequency KDE means that $\hat{V}$ rapidly and smoothly adjusts when new modes are discovered, in particular, we prove that $\hat{V}$ eventually discovers all reward modes in the CV space. More precisely,
Theorem 3.1.
where $V=V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$ is the marginal potential in the CV space if $z(x)$ is not invertible. If $z(x)$ is invertible, $V$ is the original potential $V(x)$ in the original coordinates.
The proof is in Appendix C.
Implementation details We set the potential energy beta ($\beta^{\prime}$) equal to the Langevin dynamics beta ($\beta$). This allows a physical interpretation of $\beta$: it is inversely proportional to the (unbiased) transition rates between minima of the potential. [^3] From the Kramer formula; $\text{transition rate}\propto\frac{1}{\beta}\exp(\beta\Delta V)$, where $\Delta V\propto\frac{1}{\beta^{\prime}}=\frac{1}{\beta}$. We also set the kernel width equal to the width of the Gaussian bias. This is reasonable since it is the variability of $V(x)$ that determines sensible values for both these parameters. The resulting exploration algorithm we call Adapted Metadynamics (AM), and is presented in Algorithm 1. Note that the algorithm can be extended to a batch of trajectories, where each metadynamics trajectory evolves independently, but with a shared $\hat{V}$ and $V_{\text{bias}}$ which receive updates from every trajectory in the batch. This is the version we use in our experiments - it accelerates exploration and reduces stochastic gradient noise during training.
Input : Manifold environment of terminating states $\mathcal{X}$ with reward density $r\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$. Initial state $(x_{t},p_{t})\in\mathcal{X}\times T_{x_{t}}(\mathcal{X})$. Collective variables $z=\left(z_{1},\dots,z_{k}\right)$.
Parameters : Gaussian width $\sigma=\left(\sigma_{1},\dots,\sigma_{k}\right)\in\mathbb{R}^{k}$. Gaussian height $w>0$. Stride $n\in\mathbb{Z}^{+}$. LD parameters: $\gamma,\beta$. Timestep $\Delta t$.

### v2

@@ Abstract
Generative Flow Networks (GFlowNets) are a class of generative models that sample objects in proportion to a specified reward function through a learned policy. They can be trained either on-policy or off-policy, needing a balance between exploration and exploitation for fast convergence to a target distribution. While exploration strategies for discrete GFlowNets have been studied, exploration in the continuous case remains to be investigated, despite the potential for novel exploration algorithms due to the local connectedness of continuous domains. Here, we introduce Adapted Metadynamics, a variant of metadynamics that can be applied to arbitrary black-box reward functions on continuous domains. We use Adapted Metadynamics as an exploration strategy for continuous GFlowNets. We show several continuous domains where the resulting algorithm, MetaGFN, accelerates convergence to the target distribution and discovers more distant reward modes than previous off-policy exploration strategies used for GFlowNets.
@@ 1 Introduction
Generative Flow Networks (GFlowNets) are a type of generative model that samples from a discrete space $\mathcal{X}$ by sequentially constructing objects via actions taken from a learned policy $P_{F}$ [1]. The policy $P_{F}(s,s^{\prime})$ specifies the probability of transitioning from some state $s$ to some other state $s^{\prime}$. The policy is parameterised and trained so that, at convergence, the probability of sampling an object $x\in\mathcal{X}$ is proportional to a specified reward function $R(x)$. GFlowNets offer advantages over more traditional sampling methods, such as Markov chain Monte Carlo (MCMC), by learning an amortised sampler, capable of single-shot generation of samples from the desired distribution. Since GFlowNets learn a parametric policy, they can generalise across states, resulting in higher performance across various tasks [1, 2, 3, 4, 5, 6, 7, 8, 9] and applications to conditioned molecule generation [9], maximum likelihood estimation in discrete latent variable models [7], structure learning of Bayesian networks [5], scheduling computational operations [8], and discovering reticular materials for carbon capture [10].
@@
Although originally conceived for discrete state spaces, GFlowNets have been extended to more general state spaces, such as entirely continuous spaces, or spaces that are hybrid discrete-continuous [11]. In the continuous setting, given the current state, the policy specifies a continuous probability distribution over subsequent states, and the probability density over states $x\in\mathcal{X}$ sampled with the policy is proportional to a reward density function $r(x)$. The continuous domain unlocks more applications for GFlowNets, such as molecular conformation sampling [12] and continuous control problems [13].
@@
GFlowNets are trained like reinforcement learning agents. Trajectories of states are generated either on-policy or off-policy, with the terminating state $x\in\mathcal{X}$ providing a reward signal for informing a gradient step on the policy parameters. GFlowNets therefore suffer from the same training pitfalls as reinforcement learning. One such issue is slow temporal credit assignment, which has thus far been addressed by designing more effective loss functions, such as detailed balance [14], trajectory balance [2] and sub-trajectory balance [15]. This latter approach has recently been extended by providing an energy function inductive bias at the early stages of training [16].
@@
Besides loss functions, another aspect of GFlowNet training is the exploration strategy for acquiring training samples. Exclusively on-policy learning is generally inadequate as it leads to inefficient exploration of new modes. More successful strategies therefore rely on off-policy exploration. For the discrete setting, numerous exploration strategies have been proposed including $\epsilon$-noisy with a uniform random policy, tempering, Generative Augmented Flow Networks (GAFN) [17], Thompson sampling [18] and Local Search GFlowNets [19]. While these approaches can be generalised to the continuous domain, there is limited literature benchmarking their effectiveness in this setting.
Sampling in the continuous setting is a common occurrence in various domains such as molecular modelling [20, 21] and Bayesian inference [22]. The local connectedness of a continuous domain allows for novel exploration strategies that are not directly applicable in the discrete setting. [23] compared several exploration strategies in the context of diffusion samplers, and proposed alternating between on-policy sampling and backward sampling from a fixed set of MCMC samples when training continuous GFlowNets. These samples were selected prior to training, thereby inducing a fixed pre-training cost but keeping the cost of each training trajectory constant. The authors showed effective exploration, even in high-dimensional problems. However, their MCMC approach requires access to cheap gradients of the reward landscape, which might not always be readily available [24]. In addition, there are some settings, such as in molecular conformation sampling, where MCMC approaches are known to require significantly longer timescales to overcome energy barriers than methods based on molecular dynamics (MD) simulations [25], thereby making the pre-training cost prohibitively expensive.
In this work, we propose MetaGFN, a novel exploration algorithm for continuous GFlowNets inspired by metadynamics, an enhanced sampling method widely used for molecular modelling [26]. Unlike MCMC-based sampling methods, MetaGFN operates in a general black-box setting, relying solely on an oracle for reward values, without requiring access to reward gradients. In common with standard metadynamics, MetaGFN converges quickly when a reduced-dimensional representation of the reward measure, specified by a so-called collective variable (CV) basis, is known. While our method is domain-agnostic, this feature makes MetaGFN particularly well-suited for small molecular systems, where such representations are known and often low-dimensional [27]. MetaGFN requires no pre-training and only introduces a small, constant cost for each training trajectory, which becomes negligible as the cost of reward evaluation increases.
The main contributions of this work are:
@@
  • Introducing MetaGFN, an algorithm that adapts metadynamics to black box rewards and continuous GFlowNets;
@@
  • Proving that the Adapted Metadynamics formulation underlying MetaGFN is consistent and reduces to standard metadynamics in the appropriate limit;
@@
  • Showing empirically that MetaGFN outperforms existing GFlowNets exploration strategies in various continuous environments, including alanine dipeptide conformation sampling.
@@
The rest of the paper is as follows. In Section 2 we review the theory of discrete and continuous GFlowNets as well as metadynamics and collective variables. In Section 3, we present the Adapted Metadynamics and MetaGFN algorithms. In Section 4, we evaluate MetaGFN against other approaches, showing that MetaGFN generally outperforms existing exploration strategies in various continuous environments. We finish with limitations and conclusions in Sections 5 and 6. Code for MetaGFN is available at https://github.com/dominicp6/manifold-gfn.
@@ 2.1 Discrete GFlowNets
In a GFlowNet, the network refers to a directed acyclic graph (DAG), denoted as $G=(\mathcal{S},\mathcal{A})$. Nodes represent states $s\in\mathcal{S}$, and edges represent actions $s\rightarrow s^{\prime}\in\mathcal{A}$ denoting one-way transitions between states. The DAG has two distinguishable states: a unique source state $s_{0}$, that has no incoming edges, and a unique sink state $\perp$, that has no outgoing edges.
The set of states, $\mathcal{X}\subset\mathcal{S}$, that are directly connected to the sink state are known as terminating states. GFlowNets learn forward transition probabilities, known as a forward policy $P_{F}(s^{\prime}|s)$, along the edges of the DAG so that the resulting marginal distribution over the terminal states complete trajectories (the terminal distribution) , denoted as $P^{\perp}(x)$, is proportional to a given reward function $R\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$. GFlowNets also introduce additional learnable objects, such as a backward policy $P_{B}(s|s^{\prime})$, which is a distribution over the parents of any state of the DAG, to create losses that train the forward policy. Objective functions for GFlowNets include flow matching (FM), detailed balance (DB), trajectory balance (TB) and subtrajectory balance (STB) [1, 14, 2, 15]. During training, the parameters of the flow objects are updated with stochastic gradients of the objective function applied to batches of trajectories. These trajectory batches can be obtained either directly from the current forward policy or from an alternative algorithm that encourages exploration. These approaches are known as on-policy and off-policy training respectively.
@@ 2.2 Continuous GFlowNets
Continuous GFlowNets extend the generative problem to continuous spaces [11], where the analogous quantity to the DAG is a measurable pointed graph (MPG) [28]. MPGs can model continuous spaces (e.g., Euclidean space, spheres, tori), as well as hybrid spaces, with a mix of discrete and continuous components, as often encountered in robotics, finance, and biology [29, 30, 31].
@@
    A fully-explorable space: The number of steps required to possibly reach any measurable $B\in\Sigma$ from the source state, and to guarantee to reach the sink state, with the forward reference kernel is bounded.
@@
The reward measure is a positive and finite measure $R$ over the terminating states $\mathcal{X}$. The density of $R$ is denoted as $r(x)$, for $x\in\mathcal{X}$. A flow $F$ is said to satisfy the reward-matching conditions if
@@
If a flow satisfies the reward-matching conditions and trajectories are generated by recursively sampled from the Markov kernel $P_{F}$ starting at $s_{0}$, the resulting measure over terminating states, $P^{\perp}(B)$, is proportional to the reward: $P^{\perp}(B)=\frac{R(B)}{R(\mathcal{X})}$ for any $B$ in the $\sigma$-algebra of terminating states [11].
@@ 2.3 Exploration strategies for GFlowNets
GFlowNets can reliably learn using off-policy trajectories, a key advantage over hierarchical variational models [32]. For optimal training, it is common to use a replay buffer and alternate between on-policy and off-policy (exploration) batches [33]. In the discrete case, proposed techniques to encourage exploration include $\epsilon$-noisy exploration, tempering and the incorporation of intermediate rewards [1, 17]. Exploration strategies for continuous GFlowNets are less well studied in the literature but several methods designed for discrete GFlowNets can be adapted [23]. In this work, we consider:
Local Search GFlowNets [19]: Explores by backtracking and resampling on-policy trajectories. Reconstructed trajectories with a higher reward than the original are used for training, thus encouraging the learning of high-reward modes.
Thompson sampling. [18]: Explores high-uncertainty regions by using an ensemble of policy heads with a shared torso. A random head generates the on-policy trajectory, and the loss is computed by averaging contributions over heads, where each head is independently included with probability $p$.
Noisy exploration: Explores by increasing policy uncertainty. A small constant is added to the policy variance which is gradually reduced to zero over the course of training.
Nested sampling [34]: A Markov-Chain Monte Carlo (MCMC) algorithm is used to sample from the reward distribution. Backward sampling from these terminal states are used to generate off-policy trajectories.
2.4 Metadynamics and collective Variables
Molecular dynamics (MD) simulates that dynamics of molecules using Langevin dynamics (LD) [35], a stochastic differential equation that models particle motion under friction and random noise. LD trajectories ergodically sample the molecule’s Gibbs measure, $\rho_{\beta}(x)\propto e^{-\beta V(x)}$, where $x\in\mathcal{X}$ represents atomic positions, $V(x)$ is the molecular potential, and $\beta$ is the thermodynamic beta. [^1] We review Langevin dynamics in Appendix B. However, when $V(x)$ contains multiple, deep local minima, as is common in biomolecules, then LD becomes inefficient, as it tends to get trapped in these minima, slowing exploration of the full state space.
Metadynamics overcomes this by progressively modifying the potential landscape to discourage visits to already-explored regions [26]. It achieves this by regularly depositing repulsive Gaussian biases, centered at the current state of the evolving LD trajectory. This modifies the potential to $V_{\text{total}}(x,t)=V(x)+V_{\text{bias}}(x,t)$, where $V_{\text{bias}}(x,t)$ is the cumulative bias at time $t$. Intuitively, metadynamics thus progressively transforms the landscape into more level surface, in which the system can diffuse more freely, thereby accelerating exploration (Figure 1). In the limit $t\rightarrow\infty$, the dynamics approximates free diffusion and uniformly samples the domain of $V(x)$.
Since biases are typically specified on a numerical grid, this gives rise to a memory cost that grows exponentially with dimension. Therefore, biases are typically applied along low-dimensional coordinates known as collective variables (CVs). A collective variable $z(x):\mathcal{X}\rightarrow\mathcal{Z}$ is any mapping from the original (high-dimension) state space $\mathcal{X}$ to a lower-dimensional space $\mathcal{Z}$. For potential $V(x)$ and collective variables $z$, metadynamics is guaranteed to eventually uniformly sample the domain of the marginal potential $V(z^{\prime})\vcentcolon=\int_{\mathcal{X}}\delta(z^{\prime}-z(x))V(x)\mathrm{d}x$. In molecular contexts, ideal choices of CVs correspond to the physical reaction coordinates that underly rare-event transitions (e.g. backbone dihedrals, interatomic distances) [36, 37]. In more general settings, ideal CVs should resolve rare events dynamics on $V(x)$, simplfy the landscape (retaining its essential features e.g. barriers and basins), and minimise the dimensionality of this representation. For simple systems, CVs can be derived from known symmetries or macroscopic order parameters that describe state changes. However, if the choice of CVs is not obvious, then they can learnt through data-driven methods such as Time-Lagged Independent Component Analysis (TICA) [38], manifold learning algorithms, neural network autoencoders or variational methods [39, 40, 41]. For an in-depth review on these CV-learning approaches, see [42]. Metadynamics has seen numerous extensions [43]. In the next section, we adapt the original metadynamics algorithm to the continuous black box setting. We will assume suitable CVs are given, but the theory we present is agnostic to their form–be it analytical expressions, neural networks, or non-parametric tabular mappings.
Figure 1: Illustration of metadynamics in a multi-well potential. Regular deposition of a bias leads to a total potential that gradually flattens, encouraging exploration.
@@ 3 MetaGFN: Adapted Metadynamics for GFlowNets
Metadynamics has two properties that make it well-suited as an exploration strategy. Firstly, it eventually uniformly samples the domain, ensuring exploration of all local minima. In contrast, techniques such as Local Search GFlowNets, Thompson sampling, and noisy exploration rely on localised strategies that are prone to mode locking (Section 4). Secondly, its gradual diffusion from the starting configuration allows incremental exploration, resulting in more stable training than global approaches like nested sampling. However, metadynamics requires a potential and its gradients, which are not always available in a black-box GFlowNet setting. We adapt metadynamics to this setting by interpreting the reward landscape as a potential energy surface, resulting in the exploration algorithm Adapted Metadynamics (AM) (Section 3.1) and training algorithm MetaGFN (Section 3.2).
@@
3.1 Adapted Metadynamics
Input : Manifold environment of terminating states $\mathcal{X}$ with reward density $r\vcentcolon\mathcal{X}\rightarrow\mathbb{R}$. Initial state $(x_{t},p_{t})\in\mathcal{X}\times T_{x_{t}}(\mathcal{X})$. Collective variables $z=\left(z_{1},\dots,z_{d}\right)$.
Parameters : Gaussian width $\sigma=\left(\sigma_{1},\dots,\sigma_{d}\right)\in\mathbb{R}^{d}$. Gaussian height $w>0$. Stride $n\in\mathbb{Z}^{+}$. LD parameters: $\gamma,\beta$. Timestep $\Delta t$.